For the vector space P2,find the dual basis of {1+x,1+2x,1+x+x2}?
1
Expert's answer
2020-03-19T16:56:17-0400

We are given a vector space of polynomials of degree less than or equal to 2. The given basis is, "\{1+x, 1+2x, 1+x+x^2\}" . We have to find the dual basis for it.

So, we want to find "f_1, f_2, f_3" such that, "f_i:P_2\rightarrow\mathbb{R}, i=1,2,3" are linear functionals, and,


"f_1(1+x)=1 , \\
f_2(1+2x)=1 ,\\
f_3(1+x+x^2)=1"


Consider a general element of "P_2" , say "P(x)=ax^2+bx+c, [a,b,c\in\mathbb{R}]"

As "\{1+x, 1+2x, 1+x+x^2\}" is a basis set, "P(x)" can be written uniquely as a linear combination of these 3 polynomials. Let the representation be,

"P(x)=\alpha(1+x)+\beta(1+2x)+\gamma(1+x+x^2)\\
=\gamma x^2+(\alpha+2\beta+\gamma)x+(\alpha+\beta+\gamma)"

Therefore, we must have,


"\gamma=a,\\
\alpha+2\beta+\gamma=b,\\
\alpha+\beta+\gamma=c"


Hence, "\gamma=a, \alpha=2c-b-a, \beta=b-c"

So, "P(x)" can be written as,


"P(x)=(2c-b-a)(1+x)+(b-c)(1+2x)+a(1+x+x^2)"

So, we can define "f_1, f_2, f_3" as follows:


"f_1(ax^2+bx+c)=2c-b-a\\
f_2(ax^2+bx+c)=b-c\\
f_3(ax^2+bx+c)=a"

"f_1, f_2, f_3" are linear in coefficients of the polynimial. Hence, they are linear functionals. Hence "\{f_1, f_2, f_3\}" is the dual basis for the given basis set.